OpenStudy (anonymous):

Find the number of units to be produced in order to maximize revenue if the demand function is given by p = 600x - 0.02x^2

OpenStudy (accessdenied):

This appears to be an optimization problem. We are asked to maximize the revenue, which means we need to find the revenue function and take its derivative, setting it equal to zero. To find our revenue function, think about what the demand curve represents. That is a curve representing the price of something against how many items will sell. It has the unit: money/unit sold. To find revenue then which is only in terms of money, we could multiply the information of x units sold. That gets us to the revenue function as Money/unit sold * unit sold = Money. R(x) = x * p = x (600x - 0.02x^2) If you understood everything up to this point, you should be able to proceed into the optimization step. If you need further assistance, please bump your question or tag some help!